# Directed Acyclic Outerplanar Graphs Have Constant Stack Number

### v2

@@ 3 Outerplanar DAGs have Constant Stack Number
In this section we prove Theorem 1.3, our first main result. One key ingredient is a recent result by Nöllenburg and Pupyrev [52] stating that the stack number of monotone outerplanar graphs is bounded. The idea behind our approach is to partition a given outerplanar DAG $G$ into ?transitive parts?, such that the contraction of each part into a single vertex yields a block-monotone DAG $H$. Then the result from [52] can be applied to the blocks of $H$ individually. Two things are left to do: First we show that the many stack layouts for the blocks of $H$ can be combined into a single stack layout of $H$ without requiring too many additional stacks. Then we show that each transitive part can be ?decontracted? to yield a stack layout of $G$, again without requiring too many additional stacks.
@@ 3.1 Monotone and Block-Monotone Outerplanar DAGs
Nöllenburg and Pupyrev [52] analyzed the stack number of different subclasses of outerplanar DAGs. One of their results is that monotone outerplanar DAGs with at most one vertex stacked on the base edge666 The definition of monotone in [52] allows only one vertex stacked on the base edge. This is in contrast to our definition where the base edge is (the only edge) allowed to have two vertices stacked onto it. have bounded twist number (and therefore also bounded stack number).
@@
We start with the root block, which (like all other blocks) admits an $s$-stack layout by assumption. This fulfils the invariants (I1) and (I2) trivially.
Now consider a block $B$ in level $\ell\geq 0$. Repeat the following for every cut vertex $v$ that $B$ shares with blocks $B_{1},\ldots,B_{k}$ in level $\ell+1$. For $i=1,\ldots,k$ take an $s$-stack layout of $B_{i}$ (which exists by assumption) and let $L_{i}$ and $R_{i}$ denote the sets of vertices to the left of $v$ and to the right of $v$ in the $s$-stack layout for $B_{i}$, respectively. We insert the layouts of $B_{1},\ldots,B_{k}$ directly to the left and directly to the right of $v$ such that the vertices appear in the following order (see also Figure 7):
Figure 7: Integration of the stack layouts of blocks $B_{1},\ldots,B_{k}$ around the cut vertex $v$ in the current partial stack layout. Blocks $B_{1},\ldots,B_{k}$ can use the same set of $s$ stacks except for all edges incident to $v$ which are on the $(s+1)$-th stack.
@@
Let $E^{\prime}$ denote the set of all edges that are in $B_{1},\ldots,B_{k}$. An edge from $E^{\prime}$ and an edge $e$ from a smaller level can cross only if $e$ is incident to $v$ and therefore belongs to $B$. As the edges of $B$ use at most $s+1$ stacks by invariant (I2), there are another $s+1$ stacks available for the edges in $E^{\prime}$. To assign the edges of $E^{\prime}$ to stacks we start with the $s$-stack layouts of each block and then move all edges incident to $v$ to the $(s+1)$-th stack. Observe that edges in $E^{\prime}$ belonging to different blocks can only cross if exactly one of them is incident to $v$. As the edges incident to $v$ form a star centered at $v$, we conclude that all stacks are crossing-free and $s+1$ stacks indeed suffice for $E^{\prime}$, maintaining (I2).
Finally, obverse that children of different cut vertices are separated in the layout and thus their edges do not cross. Therefore we have a $(2s+2)$-stack layout for all blocks that are already completed, maintaining (I1).
@@ 3.2 Directed $\boldsymbol{H}$-Partitions
We introduce directed $H$-partitions as a new structural tool and explore how they can be applied to reason about stack layouts of DAGs. They are not limited to DAGs or treewidth $2$ and we believe they may be applicable also for other graph classes in the future. A related tool known as layered $H$-partitions was introduced in [24] to establish that undirected planar graphs have bounded queue number and has since been used widely. Thus, we state and prove the lemmas in this section in a more general form than we actually need them. The construction of a directed $H$-partition for outerplanar DAGs is deferred to Section 3.3. The definition of directed $H$-partitions is illustrated in Figure 8.
Figure 8: A directed $H$-partition (orange) of a directed graph $G$ (left) and the quotient $H$ (right). The cut cover number of the part $P$ is $2$ as the two thick vertices cover all edges with exactly one endpoint in $P$.
@@
So each vertex $v_{P}$ in $H$ corresponds to a part in $P\in\mathcal{P}$ and a subgraph of $G$ induced by the vertices in $P$. More generally, every subgraph $B$ of $H$ corresponds to a subset $\mathcal{P}_{B}\subseteq\mathcal{P}$ of parts, and we let $G[B]$ denote the corresponding subgraph of $G$ that is induced by all vertices of $G$ contained in parts in $\mathcal{P}_{B}$.
The definition of directed $H$-partitions is very similar to the well-known concept of (undirected) $H$-partitions. The main difference and difficulty is that we need to ensure that the quotient is well-defined, i.e., the orientation of the edges between two parts is consistent. Nevertheless, many useful properties of the undirected version are inherited. In particular, if the parts are connected, then the quotient is a minor of the underlying graph, which implies that treewidth and planarity are preserved. Although not used here, we remark that the successful idea of Dujmović et al. [24] to combine $H$-partitions with layerings to so-called layered $H$-partitions is also feasible in the directed setting.
@@
Let $G$ and $H$ be DAGs and $\mathcal{P}$ be a directed $H$-partition of $G$ with cut cover number at most $w$. Further let $\operatorname{sn}(G[v_{P}])\leq s$ for each $P\in\mathcal{P}$. Then for every $h$-stack layout $\prec_{H}$ of $H$, there is a $(3wh+s)$-stack layout $\prec_{G}$ of $G$ expanding $\prec_{H}$.
@@
We expand a given $h$-stack layout $\prec_{H}$ of $H$ to a $(3wh+s)$-stack layout $\prec_{G}$ of $G$. For each part $P\in\mathcal{P}$ consider an $s$-stack layout $\prec_{P}$ of $G[v_{P}]$, which exists by assumption, and replace in $\prec_{H}$ vertex $v_{P}$ corresponding to $P$ by the vertices in $P$ ordered as in $\prec_{P}$. As in the resulting vertex ordering $\prec_{G}$ of $G$ all vertices from the same part appear consecutively (in other words, $\prec_{G}$ expands $\prec_{H}$), it follows that no two edges $e_{1},e_{2}$ in $G$ belonging to different parts in $\mathcal{P}$ cross. Therefore we may assign all edges with both endpoints in the same part to the same set of $s$ stacks.
@@
It remains to assign the edges of $G$ with endpoints in two different parts to the remaining $3wh$ stacks. For this we consider each stack $S$ in the $h$-stack layout of $H$ separately and show that all edges of $G$ corresponding to edges in $S$ can be assigned to $3w$ stacks. (An edge $vw\in E(G)$ with $v\in P$ and $w\in Q$ corresponds to the edge $v_{P}v_{Q}\in E(H)$.) First, we note that the edges in $S$ form an outerplanar subgraph of $H$. As such, it can be partitioned into three star forests [34]. Again, we can treat each star forest separately and we are left with assigning the edges in $G$ corresponding to the same star forest $F$ in $H$ to at most $w$ stacks.
@@
For this, consider two edges $e_{1},e_{2}$ of $G$ corresponding to the same star forest $F$ that cross in our chosen vertex ordering $\prec_{G}$ of $G$. If their endpoints are in four different parts in $\mathcal{P}$, then their corresponding edges in $H$ cross, which is impossible. Therefore, the endpoints of $e_{1}$ and $e_{2}$ lie in at most three different parts. Hence, there is one part containing at least two of the endpoints and we conclude that $e_{1}$ and $e_{2}$ actually correspond to the same star in $F$. Thus it suffices to consider the stars of $F$ separately and reuse the same set of $w$ stacks for all stars from $F$. Now consider all edges of $G$ corresponding to the same star $X$ of $F$. As the cut cover number of $\mathcal{P}$ is at most $w$, there is a set $V^{\prime}\subseteq V(G)$ with $|V^{\prime}|\leq w$ such that every edge of $G$ corresponding to $X$ is incident to (at least) one vertex in $V^{\prime}$. For each vertex $v^{\prime}\in V^{\prime}$ its incident edges form a star in $G$ and are therefore non-crossing in $\prec_{G}$. Thus, we can assign all incident edges at $v^{\prime}$ corresponding to $X$ to the same stack.
@@
This requires at most $w$ stacks for each star $X$ in $F$, hence also at most $w$ stacks for each star forest $F$ of $S$. To sum up, we have at most $3w$ stacks for each stack $S$ of the $h$-stack layout of $H$, and consequently at most $3w\cdot\operatorname{sn}(H)$ stacks in total.
@@
Lemma 3.8 gives a good stack layout of $G$, provided $G$ admits a directed $H$-partition with small cut cover number for some $H$ with small stack number. The notion of the cut cover number enables us to give the bound on the stack number independent of the size of the parts. We remark that without a bound on the cut cover number, there may be a twist between the vertices of two parts that is as large as the smaller of the two parts. For the next Lemma, we loosen the prerequisites by considering the blocks of $H$ separately. First, we require for each block $B$ of $H$ that the corresponding subgraph $G[B]$ has a good stack layout (for example, due to a small cut cover number of the inherited directed $H$-partition of $G[B]$). And second, the interactions between the blocks of $H$ sharing a common cut vertex $v$ are restricted.
Lemma 3.10.
Let $G$ be a DAG with a directed $H$-partition $\mathcal{P}$ such that
@@
Moreover, let $T$ be the block-cut tree of $H$ rooted at some block of $H$, such that for every cut vertex $v$ of $H$ with children blocks $C_{1},\ldots,C_{k}$ in $T$ the following holds:
  • For $i=1,\ldots,k$, the intersection $G[v]\cap G[C_{i}]$ consists of a single edge $e_{i}\in E(G[v])$.
@@
Figure 9: Left: A DAG $G$ with a directed $H$-partition as in Lemma 3.10 with $p=2$ and $t=3$. Some edge directions are omitted for better readability. Middle: $H$ and its blocks. Right: The block-cut tree of $H$.
We refer to Figure 9 for an illustration of the situation in Lemma 3.10.
Proof 3.11.
@@
We shall find a $4spt$-stack layout of $G$ with the following properties:
@@
Assuming $\prec_{G}$ satisfies (I1), the following holds:
@@
Let $T_{1}$ and $T_{2}$ be the subtrees of $T$ rooted at $B_{1}$ and $B_{2}$, respectively. First assume that $T_{1}$ and $T_{2}$ are disjoint, which in particular means that neither $B_{1}$ nor $B_{2}$ is the root of $T$. With $v_{1},v_{2}$ being the parents of $B_{1},B_{2}$ in $T$, respectively, we have $v_{1}\neq v_{2}$ since $B_{1}$ and $B_{2}$ are vertex-disjoint. Then (I1) gives that vertices of $G[B_{1}]\subseteq G[H_{v_{1}}]$ and $G[B_{2}]\subseteq G[H_{v_{2}}]$ appear in $\prec_{G}$ in disjoint intervals. Thus no edge in $G[B_{1}]$ crosses an edge in $G[B_{2}]$.
@@
If $T_{1}$ and $T_{2}$ are not vertex-disjoint, assume without loss of generality that $T_{1}\subset T_{2}$; in particular that $B_{1}$ is not the root of $T$. Then by (I1) the vertices of $G[H_{v_{1}}]$ form in $\prec_{G}$ a contiguous interval $I$. In particular, every edge in $G[B_{1}]\subseteq G[H_{v_{1}}]$ has both endpoints in $I$. As $B_{1}$ and $B_{2}$ are vertex-disjoint, we have $G[B_{2}]\cap G[H_{v_{1}}]=\emptyset$ and every edge in $G[B_{2}]$ has neither endpoint in $I$. Thus no edge in $G[B_{1}]$ crosses an edge in $G[B_{2}]$.
[…]
@@
We now construct the desired $4spt$-stack layout $\prec_{G}$ of $G$ by processing the block-cut tree $T$ from the root to the leaves. After initializing the root block, in each step we consider a cut vertex $v$ whose parent block $B$ is already processed and process all child blocks of $v$ simultaneously. In doing so, we maintain after each step (I1) and (I2) for the already processed subgraph of $G$. To initialize, the root of $T$ is a single block of $H$ and admits an $s$-stack layout expanding some vertex ordering $\prec_{H}$ of $H$ by assumption. This fulfils (I1) trivially and (I2) since the layout expands $\prec_{H}$.
[…]
@@
Figure 10: Incorporating the $s$-stack layouts of $G[X_{1}],\ldots,G[X_{4}]$ into the interval containing $G[v]$ where $G[v]\cap G[X_{i}]=\{x_{i-1},x_{i}\}$ and $Q=(x_{0},\ldots,x_{4})$ is a directed path in $G[v]$. Both, the red and the blue edges represent $s$-stack layouts of respective $G[X_{i}]$.
@@
Let $L_{i}$ and $R_{i}$ denote the sets of vertices that are to the left of $x_{i-1}$ respectively to the right of $x_{i}$ in the remaining $\prec_{i}$. Recall that there are no vertices between $x_{i-1}$ and $x_{i}$ as the vertices of $G[v]$ appear consecutively in $\prec_{i}$. We now insert $L_{i}$ immediately before $x_{i-1}$ and $R_{i}$ immediately after $x_{i}$ in the vertex ordering of the already processed graph. See also Figure 10 for a visualization. We observe that the edges of $X_{i}$ and $X_{j}$ do not cross for $|i-j|>1$. Thus, $2s$ stacks indeed suffice for all $X_{1},\ldots,X_{\ell}$ by reusing the same $s$ stacks for even indices $i$ and another $s$ stacks for odd indices $i$. As, by (I2), the vertices of $G[v]$ were consecutive in the vertex ordering before, it follows that the vertices of $G[H_{v}]$ are consecutive in the vertex ordering after those insertions, i.e., (I1) is fulfilled. Moreover, since each vertex ordering $\prec_{i}$ expands some vertex ordering of $H$, for each newly processed vertex $w$ of $H$, its corresponding subgraph $G[w]$ lies consecutively inside $L_{i}$ or inside $R_{i}$ for some $i$. Thus also (I2) is fulfilled.
{remark*}

### v3

@@ 3 Outerplanar DAGs have Constant Stack Number
In this section we prove Section 1, our first main result. One key ingredient is a recent result by Nöllenburg and Pupyrev [51] stating that the stack number of monotone outerplanar graphs is bounded. The idea behind our approach is to partition a given outerplanar DAG $G$ into “transitive parts”, such that the contraction of each part into a single vertex yields a block-monotone DAG $H$. Then the result from [51] can be applied to the blocks of $H$ individually. Two things are left to do: First we show that the many stack layouts for the blocks of $H$ can be combined into a single stack layout of $H$ without requiring too many additional stacks. Then we show that each transitive part can be “decontracted” to yield a stack layout of $G$, again without requiring too many additional stacks.
@@ 3.1 Monotone and Block-Monotone Outerplanar DAGs
Nöllenburg and Pupyrev [51] analyzed the stack number of different subclasses of outerplanar DAGs. One of their results is that monotone outerplanar DAGs with at most one vertex stacked on the base edge [^7] The definition of monotone in [51] allows only one vertex stacked on the base edge. This is in contrast to our definition where the base edge is (the only edge) allowed to have two vertices stacked onto it. have bounded twist number (and therefore also bounded stack number).
@@
We start with the root block, which (like all other blocks) admits an $s$-stack layout by assumption. This fulfills the invariants 1 and 2 trivially.
Now consider a block $B$ in level $\ell\geq 0$. We assume that $B$ is already laid out and insert its children $B_{1},\ldots,B_{k}$ into the layout. For this, repeat the following for every cut vertex $v$ that $B$ shares with blocks $B_{1},\ldots,B_{k}$ in level $\ell+1$. For $i=1,\ldots,k$ take an $s$-stack layout of $B_{i}$ (which exists by assumption) and let $L_{i}$ and $R_{i}$ denote the sets of vertices to the left of $v$ and to the right of $v$ in the $s$-stack layout for $B_{i}$, respectively. We insert the layouts of $B_{1},\ldots,B_{k}$ directly to the left and directly to the right of $v$ such that the vertices appear in the following order (see also 3.5): {sidefigure} [Uncaptioned image] Integration of the stack layouts of blocks $B_{1},\ldots,B_{k}$ around the cut vertex $v$ in the current partial stack layout. Blocks $B_{1},\ldots,B_{k}$ can use the same set of $s$ stacks except for all edges incident to $v$ which are on the $(s+1)$-th stack.
@@
Let $E^{\prime}$ denote the set of all edges that are in $B_{1},\ldots,B_{k}$. An edge from $E^{\prime}$ and an edge $e$ from a smaller level can cross only if $e$ is incident to $v$ and therefore belongs to $B$. As the edges of $B$ use at most $s+1$ stacks by invariant 2, there are another $s+1$ stacks available for the edges in $E^{\prime}$. To assign the edges of $E^{\prime}$ to stacks we start with the $s$-stack layouts of each block and then move all edges incident to $v$ to the $(s+1)$-th stack. Observe that edges in $E^{\prime}$ belonging to different blocks can only cross if exactly one of them is incident to $v$. As the edges incident to $v$ form a star centered at $v$, we conclude that all stacks are crossing-free and $s+1$ stacks indeed suffice for $E^{\prime}$, maintaining 2.
Finally, obverse that children of different cut vertices are separated in the layout and thus their edges do not cross. Therefore, we have a $(2s+2)$-stack layout for all blocks that are already completed, maintaining 1.
@@ 3.2 Directed $\boldsymbol{H}$-Partitions
We introduce directed $H$-partitions as a new structural tool and explore how they can be applied to reason about stack layouts of DAGs. They are not limited to DAGs nor to treewidth $2$, and we believe they may be applicable also for other graph classes in the future. A related tool known as layered $H$-partitions was introduced in [25] to establish that undirected planar graphs have bounded queue number and has since been used widely. Thus, we state and prove the lemmas in this section in a more general form than we actually need them. The construction of a directed $H$-partition for outerplanar DAGs is deferred to Section 3.3. The definition of directed $H$-partitions is illustrated in Figure 3.
Figure 3: A directed $H$-partition (orange) of a directed graph $G$ (left) and the quotient $H$ (right). The cut cover number of the part $P$ is $2$ as the two vertices marked in red cover all edges with exactly one endpoint in $P$.
@@
So each vertex $v_{P}$ in $H$ corresponds to a part in $P\in\mathcal{P}$ and a subgraph of $G$, denoted by $G[v_{P}]=G[P]$, that is induced by the vertices in $P$. More generally, every induced subgraph $B$ of $H$ corresponds to a subset $\mathcal{P}_{B}\subseteq\mathcal{P}$ of parts, and we let $G[B]$ denote the corresponding subgraph of $G$ that is induced by all vertices of $G$ contained in parts in $\mathcal{P}_{B}$.
The definition of directed $H$-partitions is very similar to the well-known concept of (undirected) $H$-partitions. The main difference and difficulty is that we need to ensure that the quotient is well-defined, i.e., the orientation of the edges between two parts is consistent. Nevertheless, many useful properties of the undirected version are inherited. In particular, if every part is connected, then the quotient is a minor of the underlying graph, which implies that treewidth and planarity are preserved. Although not used here, we remark that the successful idea of Dujmović et al. [25] to combine $H$-partitions with layerings to so-called layered $H$-partitions is also feasible in the directed setting.
@@
Let $G$ and $H$ be DAGs and $\mathcal{P}$ be a directed $H$-partition of $G$ with cut cover number at most $w$. Further, let $\operatorname{sn}(G[P])\leq s$ for each $P\in\mathcal{P}$. Then for every $h$-stack layout $\prec_{H}$ of $H$, there is a $(3wh+s)$-stack layout $\prec_{G}$ of $G$ expanding $\prec_{H}$.
@@
We expand a given $h$-stack layout $\prec_{H}$ of $H$ to a $(3wh+s)$-stack layout $\prec_{G}$ of $G$. For each part $P\in\mathcal{P}$ consider an $s$-stack layout $\prec_{P}$ of $G[P]$, which exists by assumption, and replace in $\prec_{H}$ vertex $v_{P}$ corresponding to $P$ by the vertices in $P$ ordered as in $\prec_{P}$. As in the resulting vertex ordering $\prec_{G}$ of $G$ all vertices from the same part appear consecutively (in other words, $\prec_{G}$ expands $\prec_{H}$), it follows that no two edges $e_{1},e_{2}$ in $G$ belonging to different parts in $\mathcal{P}$ cross. Therefore, we may assign all edges with both endpoints in the same part to the same set of $s$ stacks.
@@
It remains to assign the edges of $G$ with endpoints in two different parts to the remaining $3wh$ stacks. For this we consider each stack $S$ in the $h$-stack layout of $H$ separately and show that all edges of $G$ corresponding to edges in $S$ can be assigned to $3w$ stacks. (An edge $vw\in E(G)$ with $v\in P$ and $w\in Q$ corresponds to the edge $v_{P}v_{Q}\in E(H)$.) First, we note that the edges in $S$ form an outerplanar subgraph of $H$. As such, it can be partitioned into three star forests [34]. Again, we can treat each star forest separately, and we are left with assigning the edges in $G$ corresponding to the same star forest $F$ in $H$ to at most $w$ stacks.
@@
For this, consider two edges $e_{1},e_{2}$ of $G$ corresponding to the same star forest $F$ that cross in our chosen vertex ordering $\prec_{G}$ of $G$. If their endpoints are in four different parts in $\mathcal{P}$, then their corresponding edges in $H$ cross, which is impossible. Therefore, the endpoints of $e_{1}$ and $e_{2}$ lie in at most three different parts. Hence, there is one part containing at least two of the endpoints, and we conclude that $e_{1}$ and $e_{2}$ actually correspond to the same star in $F$. Thus it suffices to consider the stars of $F$ separately and reuse the same set of $w$ stacks for all stars from $F$. Now consider all edges of $G$ corresponding to the same star $X$ of $F$. As the cut cover number of $\mathcal{P}$ is at most $w$, there is a set $V^{\prime}\subseteq V(G)$ with $|V^{\prime}|\leq w$ such that every edge of $G$ corresponding to $X$ is incident to (at least) one vertex in $V^{\prime}$. For each vertex $v^{\prime}\in V^{\prime}$ its incident edges form a star in $G$ and are therefore non-crossing in $\prec_{G}$. Thus, we can assign all incident edges at $v^{\prime}$ corresponding to $X$ to the same stack.
@@
This requires at most $w$ stacks for each star $X$ in $F$, hence also at most $w$ stacks for each star forest $F$ of $S$. To sum up, we have at most $3w$ stacks for each stack $S$ of the $h$-stack layout of $H$. Including the $s$ stacks from the beginning, this yields at most $3wh+s$ stacks in total.
@@
Lemma 3.8 gives a good stack layout of $G$, provided $G$ admits a directed $H$-partition with small cut cover number for some $H$ with small stack number. The notion of the cut cover number enables us to give the bound on the stack number independently of the size of the parts. We remark that without a bound on the cut cover number, there may be a twist between the vertices of two parts that is as large as the smaller of the two parts. For the next Section, we loosen the prerequisites by considering the blocks of $H$ separately. First, we require for each block $B$ of $H$ that the corresponding subgraph $G[B]$ has a good stack layout (for example, due to a small cut cover number of the inherited directed $H$-partition of $G[B]$). And second, the interactions between the blocks of $H$ sharing a common cut vertex $v$ are restricted.
{restatable}
lemmaHPartitionToStackLayout Let $G$ be a DAG with a directed $H$-partition $\mathcal{P}$ such that
@@
Moreover, let $T$ be the block-cut tree of $H$ rooted at some block of $H$, such that for every cut vertex $v$ of $H$ with child blocks $C_{1},\ldots,C_{k}$ in $T$ the following holds:
  • For $i=1,\ldots,k$, the intersection of $G[v]$ with the neighborhood of $G[C_{i}-v]$ consists of a single edge $e_{i}\in E(G[v])$.
@@
Figure 4: Left: A DAG $G$ with a directed $H$-partition as in 3.9 with $p=2$ and $t=3$. Some edge directions are omitted for better readability. Middle: $H$ and its blocks. Right: The block-cut tree of $H$.
We refer to Figure 4 for an illustration of the situation in 3.9. In the following proof, we denote the neighborhood of a subgraph $G^{\prime}$ of some graph $G$ by $N(G^{\prime})$.
Proof 3.10.
@@
We shall find a $4spt$-stack layout of $G$ whose vertex ordering $\prec_{G}$ has the following properties:
@@
Assuming $\prec_{G}$ satisfies 1, the following holds:
@@
Let $T_{1}$ and $T_{2}$ be the subtrees of $T$ rooted at $B_{1}$ and $B_{2}$, respectively. First assume that $T_{1}$ and $T_{2}$ are disjoint, which in particular means that neither $B_{1}$ nor $B_{2}$ is the root of $T$. With $v_{1},v_{2}$ being the parents of $B_{1},B_{2}$ in $T$, respectively, we have $v_{1}\neq v_{2}$ since $B_{1}$ and $B_{2}$ are vertex-disjoint. Then 1 gives that vertices of $G[B_{1}]\subseteq G[H_{v_{1}}]$ and $G[B_{2}]\subseteq G[H_{v_{2}}]$ appear in $\prec_{G}$ in disjoint intervals. Thus no edge in $G[B_{1}]$ crosses an edge in $G[B_{2}]$.
@@
If $T_{1}$ and $T_{2}$ are not vertex-disjoint, assume without loss of generality that $T_{1}\subset T_{2}$; in particular that $B_{1}$ is not the root of $T$. Then by 1 the vertices of $G[H_{v_{1}}]$ form in $\prec_{G}$ a contiguous interval $I$. In particular, every edge in $G[B_{1}]\subseteq G[H_{v_{1}}]$ has both endpoints in $I$. As $B_{1}$ and $B_{2}$ are vertex-disjoint, we have $G[B_{2}]\cap G[H_{v_{1}}]=\emptyset$ and every edge in $G[B_{2}]$ has neither endpoint in $I$. Thus no edge in $G[B_{1}]$ crosses an edge in $G[B_{2}]$.
Claim 1 allows us to reuse the same set of stacks for vertex-disjoint blocks. With this in mind, we partition the blocks of $H$ into two sets $\mathcal{B}_{\text{odd}}$ and $\mathcal{B}_{\text{even}}$ containing the blocks with an odd, respectively even number of cut vertices on their path to the root in $T$. Then it is enough to use a set of $2spt$ stacks for blocks in $\mathcal{B}_{\text{odd}}$ and a set of $2spt$ different stacks for $\mathcal{B}_{\text{even}}$, giving the desired $4spt$ stacks in total. Observe that within the same set of blocks, say $\mathcal{B}_{\text{odd}}$, two blocks are either again vertex-disjoint (and thus non-crossing by Claim 1) or have a common parent in $T$. Thus it is left to consider a single cut vertex and its child blocks.
@@
We now construct the desired $4spt$-stack layout $\prec_{G}$ of $G$ by processing the block-cut tree $T$ from the root to the leaves. After initializing the root block, in each step we consider a cut vertex $v$ whose parent block $B$ is already processed and process all child blocks of $v$ simultaneously. In doing so, we maintain after each step 1 and 2 for the already processed subgraph of $G$. To initialize, the root of $T$ is a single block of $H$ and admits an $s$-stack layout expanding some vertex ordering $\prec_{H}$ of $H$ by assumption. This fulfills 1 trivially and 2 since the layout expands $\prec_{H}$.
Now for a step, consider a cut vertex $v$ whose parent block is already processed and let $C_{1},\ldots,C_{k}$ be the child blocks of $v$ in $T$. By the assumptions of the Section, for each $i=1,\ldots,k$ the intersection $G[v]\cap N(G[C_{i}-v])$ consists of a single edge $e_{i}\in E(G[v])$, the edges $e_{1},\ldots,e_{k}$ are covered with at most $p$ directed paths $Q_{1},\ldots,Q_{p}$ in $G[v]$, and for each edge $e\in E(G[v])$ we have $e=e_{i}$ for at most $t$ indices $i\in\{1,\ldots,k\}$. Since all of $C_{1},\ldots,C_{k}$ are in $\mathcal{B}_{\text{even}}$ or all in $\mathcal{B}_{\text{odd}}$, we have a set of $2spt$ stacks at our disposal. Reserve $p$ pairwise disjoint sets of stacks of size $2st$, one per path. For a fixed path $Q$, group these $2st$ stacks further into $t$ disjoint subsets of size $2s$, such that for each subset each edge $e$ of $Q$ corresponds to at most one index $i\in\{1,\ldots,k\}$ with $e=e_{i}$. It is left to show that we can find a $2s$-stack layout for a fixed path $Q=(x_{0},\ldots,x_{\ell})$ of length $\ell$ and a set of blocks $X_{1},\ldots,X_{\ell}$, where $G[v]\cap N(G[X_{j}-v])$ is exactly the edge $x_{j-1}x_{j}$ with $j=1,\dots,\ell$ (having one block per edge is the most difficult case, having less only makes it easier), refer to Figure 5. Note that each child $C_{i}$ is dealt with: It is either in $\mathcal{B}_{\text{even}}$ or $\mathcal{B}_{\text{odd}}$, it is attached to one of the $p$ paths, and it is contained in one of the $t$ subsets for the path within which $C_{i}$ does not share the edge $e_{i}$ with some other child. That is, each of the $X_{j}$’s takes care of at most one $C_{i}$.
Figure 5: The situation of 3.9 (see Figure 4, in particular the three blocks $C_{1},C_{2},C_{3}$ attached to the same edge in $G[v]$) simplified in two ways: We only consider one directed path $Q\subseteq G[v]$ and we have exactly one block $X_{j}$ at each edge $x_{j-1}x_{j}$, instead of at most $t$ (which is paid for by a factor of $t$) or none (for which the respective $X_{j}$ can simply be ignored). That is, each block $X_{j}$ represents at most one child block $C_{i}$ of $v$, where $e_{i}=x_{j-1}x_{j}$. Note that the colors match those used in the layout in 3.10.
We are now in the situation illustrated in Figure 5 and by the assumptions of the Section, for each block $X_{j}$ the corresponding $G[X_{j}]$ admits an $s$-stack layout $\prec_{j}$ expanding some vertex ordering of $H$. For every vertex $w$ in $X_{j}$ the vertices of $G[w]$ appear consecutively in $\prec_{j}$, and this holds in particular for $w=v$. We remove from $\prec_{j}$ all vertices in $G[v]$ except for $x_{j-1}$ and $x_{j}$. As $G[v]$ and $G[X_{j}-v]$ are connected only via $x_{j-1}$ and $x_{j}$, this does not remove any edge of $G$ corresponding to an edge in $X_{j}-v$.
{cfigure}[Uncaptioned image]
@@
Incorporating the $s$-stack layouts of $G[X_{1}],\ldots,G[X_{4}]$ (compare Figure 5) into the interval containing $G[v]$, where $G[v]\cap N(G[X_{i}-v])=\{x_{i-1},x_{i}\}$ and $Q=(x_{0},\ldots,x_{4})$ is a directed path in $G[v]$. Both, the red and the blue edges represent $s$-stack layouts of respective $G[X_{i}]$.
@@
Let $L_{j}$ and $R_{j}$ denote the sets of vertices that are to the left of $x_{j-1}$ respectively to the right of $x_{j}$ in the remaining $\prec_{j}$. Recall that there are no vertices between $x_{j-1}$ and $x_{j}$ as the vertices of $G[v]$ appear consecutively in $\prec_{j}$. We now insert $L_{j}$ immediately before $x_{j-1}$ and $R_{j}$ immediately after $x_{j}$ in the vertex ordering of the already processed graph. See also 3.10 for a visualization. We observe that the edges of $X_{j}$ and $X_{j^{\prime}}$ do not cross for $|j-j^{\prime}|>1$. Thus, $2s$ stacks indeed suffice for all $X_{1},\ldots,X_{\ell}$ by reusing the same $s$ stacks for even indices $j$ and another $s$ stacks for odd indices $j$.
To finish the proof, recall that we maintain 1 and 2 after each step for the already processed subgraph of $G$. For this, note that by processing its child blocks, $v$ is turned from a non-cut vertex to a cut vertex, and thus we may assume 2 but need to satisfy 1 for $v$. As, by 2, the vertices of $G[v]$ were consecutive in the vertex ordering before, it follows that the vertices of $G[H_{v}]$ are consecutive in the vertex ordering after those insertions, i.e., 1 is fulfilled. Moreover, since each vertex ordering $\prec_{j}$ expands some vertex ordering of $H$, we conclude that for each newly processed vertex $w$ of $H$, its corresponding subgraph $G[w]$ lies consecutively inside $L_{j}$ or inside $R_{j}$ for some $j$. Thus also 2 is fulfilled.
Remark 3.11.